OpenStudy (anonymous):

The lengths of the legs of a right triangle are 6 inches and 8 inches. What is the length of the hypotenuse?

OpenStudy (calculusxy):

Use the Pythagorean theorem

OpenStudy (calculusxy):

a^2 + b^2 = c^2

OpenStudy (acxbox22):

6^2+8^2=100 \[\sqrt{100}=10\]

OpenStudy (calculusxy):

A and b can be used for the legs and c will be the hypotenuse

OpenStudy (anonymous):

We have gone over this before haven't we? >.<